OpenStudy (anonymous):

Write the equation of a circle with a center at (6,7) and a diameter of 4?

OpenStudy (anonymous):

\[(x-6)^2+(y-7)^2=16\]

OpenStudy (anonymous):

Awh dang you're good! lol :D thanks buddy!!

OpenStudy (anonymous):

your welcome mate ;)

OpenStudy (anonymous):

lol :D

OpenStudy (hba):

\[\huge\ (x-a)^2+(y-b)^2=r^2\]

OpenStudy (hba):

@itsmylife Giving direct answers is not allowed on Os. Please check the CoC- http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

The standard equation for a circle is: (x - a)² + (y-b)² = r² Just plug in the values... the center of the circle is (a, b) in the equation. The diameter divided by 2 is the radius..

OpenStudy (anonymous):

thanks guys!! And @hba people do it all the time, I do not but people do it alot!!

OpenStudy (hba):

Wrong is Wrong,Either you do it all the time or once.

OpenStudy (anonymous):

@hba are ya running this website? second question ya just wrote formula of circle in bigger font ya think your very smart lol learn before ya teach , check your self before licking shoe of mods i don give a damn to your stupid attitude , ya can get me a warning again ma apology but i aint giving answer n ya do it too, mind yor business

OpenStudy (anonymous):

@hba if yor thinking to be a mod lol that would be your dream , they make mod to someone whos got sense to think not to sillies

OpenStudy (anonymous):

ok guys! He helped me and I got the things I needed!! It's fine. And yeah it's wrong, but I don't see you telling other people this @hba so I'm sorry.. People will explain it and if you would've gave him a minute he would've so i'm sorry!!!!

OpenStudy (anonymous):

anything else im here for ya

OpenStudy (anonymous):

@BrittanyFrench ya dont need to be sorry , are ya crazy or wat, what are ya sorry for, all the kids do that here i aint scared of anything

OpenStudy (anonymous):

@itsmylife I will email you a very angry message lol. Chill out please!!! It's fine..

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@ashlynn143 thanks! :D I may need help more lol

OpenStudy (anonymous):

anything else

OpenStudy (anonymous):

Yeah maybe, I'm gonna close this one and open another one lol

OpenStudy (anonymous):

ok